# Convergence to equilibrium for a degenerate three species reaction-diffusion system

### v1

@@ 2. The case of $d_{b}=0$
The idea is to connect entropy dissipation with the missing $\delta_{B}$ term in (12) so that we can apply Grönwall inequality to have a sub-exponential decay. We begin by proving an uniform integrability estimate for $a$ and $c$ in a parabolic cylinder of unit height. A similar estimate was obtained for a degenerate four species model in [EMT20, Lemma 3.12, p.4343].
@@
$\partial_{t}\left(\phi_{\tau}a\right)-d\Delta\left(\phi_{\tau}a\right)&=a\partial_{t}\phi_{\tau}+\phi_{\tau}(c-ab+(d_{a}-d)\Delta a)\qquad\mbox{ in }\Omega_{\tau,\tau+2}\\ \phi_{\tau}\nabla a\cdot n&=0\qquad\qquad\qquad\qquad\ \qquad\qquad\qquad\ \mbox{ on }\partial\Omega_{\tau,\tau+2}\\ \phi_{\tau}(\tau)a(\tau,x)&=0\qquad\qquad\qquad\qquad\qquad\qquad\qquad\ \ \ \mbox{ in }\Omega\end{aligned}\right.$
@@
(14) $\partial_{t_{1}}(\phi_{\tau}(t_{1}+\tau)a(t_{1}+\tau,x))&-d\Delta(\phi_{\tau}(t_{1}+\tau)a(t_{1}+\tau,x))\\ =a(t_{1}+\tau,x)\partial_{t_{1}}\phi_{\tau}(t_{1}&+\tau)+\phi_{\tau}(t_{1}+\tau)(c-ab+(d_{a}-d)\Delta a)\qquad\mbox{ in }\Omega_{0,2}\\ \phi_{\tau}(t_{1}+\tau)\nabla a(t_{1}+\tau,x)\cdot n&=0\qquad\qquad\quad\qquad\qquad\qquad\qquad\qquad\qquad\mbox{ on }\partial\Omega_{0,2}\\ \phi_{\tau}(0+\tau)a(0+\tau,x)&=0\qquad\qquad\qquad\qquad\quad\qquad\qquad\qquad\qquad\mbox{ in }\Omega\end{aligned}\right.$
@@
(15) $\partial_{t_{1}}(\phi_{\tau}(t_{1}+\tau)c(t_{1}+\tau,x))&-d\Delta(\phi_{\tau}(t_{1}+\tau)c(t_{1}+\tau,x))\\ =c(t_{1}+\tau,x)\partial_{t_{1}}\phi_{\tau}(t_{1}&+\tau)+\phi_{\tau}(t_{1}+\tau)(ab-c+(d_{c}-d)\Delta c)\qquad\mbox{ in }\Omega_{0,2}\\ \phi_{\tau}(t_{1}+\tau)\nabla c(t_{1}+\tau,x)\cdot n&=0\qquad\quad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\mbox{ on }\partial\Omega_{0,2}\\ \phi_{\tau}(0+\tau)c(0+\tau,x)&=0\qquad\qquad\qquad\qquad\quad\qquad\qquad\qquad\qquad\mbox{ in }\Omega\end{aligned}\right.$
@@
(18) $\partial_{t_{1}}\psi(t_{1},x)+d\Delta\psi(t_{1},x)&=-\theta(t_{1},x)\qquad\mbox{ for }(t_{1},x)\in\Omega_{0,2}\\ \nabla\psi(t_{1},x)\cdot n&=0\qquad\ \ \qquad\mbox{ for }(t_{1},x)\in(0,2)\times\partial\Omega\\ \psi(2,x)&=0\qquad\qquad\mbox{ for }x\in\Omega\end{aligned}\right.$
@@
(19) $\|\Delta\psi\|_{\mathrm{L}^{p^{\prime}}(\Omega_{0,2})}\leq\frac{2C_{SOR}(\Omega,N,p^{\prime})}{d_{a}+d_{c}}\|\theta\|_{\mathrm{L}^{p^{\prime}}(\Omega_{0,2})}$
@@
$+\int_{0}^{2}\int_{\Omega}\left(\int_{0}^{t_{1}}\int_{\Omega}G_{d}(t_{1},s,x,y)\phi_{\tau}(s+\tau)(c-ab+(d_{a}-d)\Delta a)(s,y)\,{\rm d}y\,{\rm d}s\right)\theta(t_{1},x)\,{\rm d}x\,{\rm d}t_{1}$
@@
The following Green’s function estimate is available from [Mor83][CK15]: there exists a constant $K_{1}>0$, depending only on the domain, such that
@@
$\left\|g_{d}\right\|_{L^{1+\frac{2}{N}}(\Omega_{0,2})}\leq K_{2}$
for some constant $K_{2}>0$. Observe that there exists a $q<p$ such that
$1+\frac{1}{p}=\frac{1}{1+\frac{2}{N}}+\frac{1}{q}.$
@@
$\left\|\int_{0}^{t_{1}}\int_{\Omega}G_{d}(t_{1},s,x,y)a(s+\tau,y)\partial_{s}\phi_{\tau}(s+\tau)\,{\rm d}y\,{\rm d}s\right\|_{\mathrm{L}^{p}(\Omega_{0,2})}$ $\leq\left\|g_{d}\right\|_{\mathrm{L}^{1+\frac{2}{n}}(\Omega_{0,2})}\left\|a(\cdot+\tau,\cdot)\partial_{s}\phi_{\tau}(\cdot+\tau)\right\|_{\mathrm{L}^{q}(\Omega_{0,2})}$
$\leq K_{2}\left\|a(\cdot+\tau,\cdot)\partial_{s}\phi_{\tau}(\cdot+\tau)\right\|_{\mathrm{L}^{q}(\Omega_{0,2})}$
@@
$+C_{SOR}(\Omega,N,p^{\prime})\frac{|d_{a}-d_{c}|}{d_{a}+d_{c}}\left\|\phi_{\tau}(\cdot+\tau)a(\cdot+\tau,\cdot)\right\|_{\mathrm{L}^{p}(\Omega_{0,2})}\|\theta\|_{\mathrm{L}^{p^{\prime}}(\Omega_{0,2})}$
@@
$+C_{SOR}(\Omega,N,p^{\prime})\frac{|d_{a}-d_{c}|}{d_{a}+d_{c}}\left\|\phi_{\tau}(\cdot+\tau)c(\cdot+\tau,\cdot)\right\|_{\mathrm{L}^{p}(\Omega_{0,2})}\|\theta\|_{\mathrm{L}^{p^{\prime}}(\Omega_{0,2})}$
@@
$+C_{SOR}(\Omega,N,p^{\prime})\frac{|d_{a}-d_{c}|}{d_{a}+d_{c}}\left\|\phi_{\tau}(\cdot+\tau)\left(a+c\right)(\cdot+\tau,\cdot)\right\|_{\mathrm{L}^{p}(\Omega_{0,2})}\|\theta\|_{\mathrm{L}^{p^{\prime}}(\Omega_{0,2})}$
@@
(23) $\left\|\phi_{\tau}(\cdot+\tau)(a+c)(\cdot+\tau,\cdot)\right\|_{\mathrm{L}^{p}(\Omega_{0,2})}\leq\frac{K_{2}M^{\alpha}M_{1}^{\alpha}|\Omega|^{\alpha}}{\left(1-C_{SOR}(\Omega,N,p^{\prime})\frac{|d_{a}-d_{c}|}{d_{a}+d_{c}}\right)}\left\|\left(a+c\right)(\cdot+\tau,\cdot)\partial_{s}\phi_{\tau}(\cdot+\tau)\right\|_{\mathrm{L}^{p}(\Omega_{0,2})}^{1-\alpha}$
@@
As the diffusion coefficients $d_{a}$ and $d_{c}$ satisfy the closeness condition (10), employing the $p^{\textrm{th}}$ order integrability estimate [CDF14, Proposition 1.1, p.1186] (see Theorem 12 from the Appendix for the precise statement), we arrive at
$\beta_{0}=\|a+c\|_{\mathrm{L}^{p}(\Omega_{0,1})}\leq\left(1+\max\{d_{a},d_{c}\}\frac{|d_{c}-d_{a}|\,C^{PRC}_{\frac{d_{a}+d_{c}}{2},p^{\prime}}}{2-|d_{c}-d_{a}|\,C^{PRC}_{\frac{d_{a}+d_{c}}{2},p^{\prime}}}\right)\|a_{0}+c_{0}\|_{\mathrm{L}^{p}(\Omega)}$
@@
$\left(b(t,x)\right)^{N}\leq 2^{N-1}\left(\left(b_{0}(x)\right)^{N}+\left(\int_{0}^{t}c(s,x)\,{\rm d}s\right)^{N}\right)\leq 2^{N-1}(1+t)^{N-1}\left(\left(b_{0}(x)\right)^{N}+\int_{0}^{t}\left(c(s,x)\right)^{N}\,{\rm d}s\right)$
@@
(24) $\left\|b(t,\cdot)\right\|_{\mathrm{L}^{N}(\Omega)}^{N}\leq 2^{N-1}(1+t)^{N-1}\Big{(}\left\|b_{0}\right\|_{\mathrm{L}^{N}(\Omega)}^{N}+\left\|c\right\|_{\mathrm{L}^{N}(\Omega_{t})}^{N}\Big{)}$
@@
$\left\|c\right\|_{\mathrm{L}^{N}(\Omega_{\tau,\tau+1})}\leq\left\|c\right\|_{\mathrm{L}^{p}(\Omega_{\tau,\tau+1})}\,\left|\Omega\right|^{\frac{p-N}{p}}$
@@
$\left\|b(t,\cdot)\right\|_{\mathrm{L}^{N}(\Omega)}^{N}\leq 2^{N-1}(1+t)^{N-1}\left(\left\|b_{0}\right\|_{\mathrm{L}^{\infty}(\Omega)}^{N}\,\left|\Omega\right|+C_{0}^{N}\,\left|\Omega\right|^{\frac{N(p-N)}{p}}\,(1+t)\right)$
@@
$K_{3}^{\frac{N(N-1)}{N-2}}:=2^{N-1}\left(M_{2}|\Omega|\right)^{\frac{N}{N-2}}\left(\left\|b_{0}\right\|_{\mathrm{L}^{\infty}(\Omega)}^{N}\,\left|\Omega\right|+C_{0}^{N}\,\left|\Omega\right|^{\frac{N(p-N)}{p}}\right)$
@@
$\left\|b(t,\cdot)\right\|_{\mathrm{L}^{N}(\Omega)}\leq K_{3}^{\frac{N-1}{N-2}}\left(M_{2}|\Omega|\right)^{-\frac{1}{N-2}}(1+t)$
@@
$\left\|b(t,\cdot)\right\|_{\mathrm{L}^{\frac{N}{2}}(\Omega)}\leq\left\|b(t,\cdot)\right\|_{\mathrm{L}^{N}(\Omega)}^{\frac{N-2}{N-1}}\,\left\|b(t,\cdot)\right\|_{\mathrm{L}^{1}(\Omega)}^{\frac{1}{N-1}}$
@@
$\left\|c\right\|_{\mathrm{L}^{\frac{7}{2}}(\Omega_{\tau,\tau+1})}$ $\leq K_{c}(1+t)^{\mu_{c}}\qquad\mbox{ for }\,\tau\geq 0,$
@@
$\frac{1}{p}=\frac{j}{N}+\big{(}\frac{1}{r}-\frac{m}{N}\big{)}\alpha\ +\frac{1-\alpha}{q}\qquad\mbox{ and }\qquad\frac{j}{N}\leq\alpha\leq 1.$
@@
$\left\|\sqrt{a}(t,\cdot)\right\|_{\mathrm{L}^{\frac{10}{3}}(\Omega)}\leq K_{5}\left\|\nabla\sqrt{a}(t,\cdot)\right\|_{\mathrm{L}^{2}(\Omega)}^{\frac{3}{5}}\,\left\|\sqrt{a}(t,\cdot)\right\|_{\mathrm{L}^{2}(\Omega)}^{\frac{2}{5}}+K_{6}\left\|\sqrt{a}(t,\cdot)\right\|_{\mathrm{L}^{2}(\Omega)}$
@@
$=2^{\frac{7}{3}}K_{5}^{\frac{10}{3}}\left\|\nabla\sqrt{a}(t,\cdot)\right\|_{\mathrm{L}^{2}(\Omega)}^{2}\,M_{1}^{\frac{4}{3}}\,\left|\Omega\right|^{\frac{4}{3}}+2^{\frac{7}{3}}K_{6}^{\frac{10}{3}}\,M_{1}^{\frac{10}{3}}\,\left|\Omega\right|^{\frac{10}{3}}$
@@
$\leq 2^{\frac{2}{3}}(1+t)^{\frac{5}{3}}\left(\left\|b_{0}\right\|_{\mathrm{L}^{\infty}(\Omega)}^{\frac{5}{3}}\,\left|\Omega\right|+K_{7}^{\frac{5}{3}}\right)$
@@
As $a$ is a positive subsolution of the parabolic equation $\partial_{t}a-d_{a}\Delta a\leq c$, integrability estimation (11) yields
$\|a\|_{L^{4.5}(\Omega_{0,2})}\leq C_{IE}(\Omega,d_{a})\|c\|_{L^{\frac{5}{3}}(\Omega_{0,2})}\leq C_{IE}(\Omega,d_{a})K_{7}.$
@@
$-[\partial_{t}\phi+d_{c}\Delta\phi]=$ $\Theta\qquad\ \ \ \Omega_{0,2}$
$\nabla\phi.n=$ $0\qquad(0,2)\times\partial\Omega$
ϕ(2)=italic-$\phi(2)=$ $0\qquad\ \ \ \ \Omega$
We have $\phi\geq 0$, and the following estimates for a constant $C_{q,d_{c}}>0$( $q\in(1,\infty)$, arbitrary)[Pie10][Ama85]
$\|\phi_{t}\|_{L^{q}(\Omega_{0,2})}+\|\Delta\phi\|_{L^{q}(\Omega_{0,2})}+\sup\limits_{s\in[0,2]}\|\phi(s)\|_{L^{q}(\Omega)}+\|\phi\|_{L^{q}((0,2)\times\partial\Omega)}\leq C_{q,d_{c}}\|\Theta\|_{L^{q}(\Omega_{0,2})}.$

### v2

@@ 2. The case of $d_{b}=0$
The idea is to connect entropy dissipation with the missing $\|\delta_{B}\|^{2}_{\mathrm{L}^{2}(\Omega)}$ term in (12) so that we can apply Grönwall inequality to have a sub-exponential decay. We begin by proving an uniform integrability estimate for $a$ and $c$ in a parabolic cylinder of unit height. A similar estimate was obtained for a degenerate four species model in [7, Lemma 3.12, p.4343].
@@
$\partial_{t}\left(\phi_{\tau}a\right)-d\Delta\left(\phi_{\tau}a\right)&=a\partial_{t}\phi_{\tau}+\phi_{\tau}(c-ab+(d_{a}-d)\Delta a)\qquad\mbox{ in }\Omega_{\tau,\tau+2}\\ \phi_{\tau}\nabla a\cdot n&=0\qquad\qquad\qquad\qquad\ \qquad\qquad\qquad\ \ \ \ \mbox{ on }\partial\Omega_{\tau,\tau+2}\\ \phi_{\tau}(\tau)a(\tau,x)&=0\qquad\qquad\qquad\qquad\qquad\qquad\qquad\ \ \ \ \ \mbox{ in }\Omega.\end{aligned}\right.$
@@
(17) $\partial_{t_{1}}(\phi_{\tau}(t_{1}+\tau)a(t_{1}+\tau,x))&-d\Delta(\phi_{\tau}(t_{1}+\tau)a(t_{1}+\tau,x))\\ =a(t_{1}+\tau,x)\partial_{t_{1}}\phi_{\tau}(t_{1}&+\tau)+\phi_{\tau}(t_{1}+\tau)(c-ab+(d_{a}-d)\Delta a)\qquad\mbox{ in }\Omega_{0,2}\\ \phi_{\tau}(t_{1}+\tau)\nabla a(t_{1}+\tau,x)\cdot n&=0\qquad\qquad\quad\qquad\qquad\ \qquad\qquad\qquad\qquad\mbox{ on }\partial\Omega_{0,2}\\ \phi_{\tau}(0+\tau)a(0+\tau,x)&=0\qquad\qquad\qquad\qquad\quad\qquad\qquad\qquad\qquad\ \ \mbox{ in }\Omega.\end{aligned}\right.$
@@
(18) $\partial_{t_{1}}(\phi_{\tau}(t_{1}+\tau)c(t_{1}+\tau,x))&-d\Delta(\phi_{\tau}(t_{1}+\tau)c(t_{1}+\tau,x))\\ =c(t_{1}+\tau,x)\partial_{t_{1}}\phi_{\tau}(t_{1}&+\tau)+\phi_{\tau}(t_{1}+\tau)(ab-c+(d_{c}-d)\Delta c)\qquad\mbox{ in }\Omega_{0,2}\\ \phi_{\tau}(t_{1}+\tau)\nabla c(t_{1}+\tau,x)\cdot n&=0\qquad\quad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\ \mbox{ on }\partial\Omega_{0,2}\\ \phi_{\tau}(0+\tau)c(0+\tau,x)&=0\qquad\qquad\qquad\qquad\quad\qquad\qquad\qquad\qquad\ \ \mbox{ in }\Omega.\end{aligned}\right.$
@@
(21) $\partial_{t_{1}}\psi(t_{1},x)+d\Delta\psi(t_{1},x)&=-\theta(t_{1},x)\qquad\mbox{in}\ \Omega_{0,2}\\ \nabla\psi(t_{1},x)\cdot n&=0\qquad\ \ \qquad\ \ \mbox{on}\ (0,2)\times\partial\Omega\\ \psi(2,x)&=0\qquad\qquad\quad\ \mbox{in}\ \Omega.\end{aligned}\right.$
@@
(22) $\|\Delta\psi\|_{\mathrm{L}^{p^{\prime}}(\Omega_{0,2})}\leq\frac{2C_{SOR}(\Omega,N,p^{\prime})}{d_{a}+d_{c}}\|\theta\|_{\mathrm{L}^{p^{\prime}}(\Omega_{0,2})}.$
@@
$+\int_{0}^{2}\int_{\Omega}\left(\int_{0}^{t_{1}}\int_{\Omega}G_{d}(t_{1},s,x,y)\phi_{\tau}(s+\tau)(c-ab+(d_{a}-d)\Delta a)(s,y)\,{\rm d}y\,{\rm d}s\right)\theta(t_{1},x)\,{\rm d}x\,{\rm d}t_{1}.$
@@
The following Green’s function estimate is available from [10][3]: there exists a constant $K_{1}>0$, depending only on the domain, such that
@@
$\left\|g_{d}\right\|_{\mathrm{L}^{z}(\Omega_{0,2})}\leq K_{2}\qquad\forall z\in\left[1,1+\frac{N}{2}\right)$
for some constant $K_{2}>0$ depending on $z$. We, in particular choose ${z=1+\frac{1}{N}}$. Observe that there exists a $q<p$ such that
$1+\frac{1}{p}=\frac{1}{1+\frac{1}{N}}+\frac{1}{q}.$
@@
$\left\|\int_{0}^{t_{1}}\int_{\Omega}G_{d}(t_{1},s,x,y)a(s+\tau,y)\partial_{s}\phi_{\tau}(s+\tau)\,{\rm d}y\,{\rm d}s\right\|_{\mathrm{L}^{p}(\Omega_{0,2})}$ $\leq\left\|g_{d}\right\|_{\mathrm{L}^{1+\frac{1}{n}}(\Omega_{0,2})}\left\|a(\cdot+\tau,\cdot)\partial_{s}\phi_{\tau}(\cdot+\tau)\right\|_{\mathrm{L}^{q}(\Omega_{0,2})}$
$\leq K_{2}\left\|a(\cdot+\tau,\cdot)\partial_{s}\phi_{\tau}(\cdot+\tau)\right\|_{\mathrm{L}^{q}(\Omega_{0,2})}.$
@@
$+C_{SOR}(\Omega,N,p^{\prime})\frac{|d_{a}-d_{c}|}{d_{a}+d_{c}}\left\|\phi_{\tau}(\cdot+\tau)a(\cdot+\tau,\cdot)\right\|_{\mathrm{L}^{p}(\Omega_{0,2})}\|\theta\|_{\mathrm{L}^{p^{\prime}}(\Omega_{0,2})}.$
@@
$+C_{SOR}(\Omega,N,p^{\prime})\frac{|d_{a}-d_{c}|}{d_{a}+d_{c}}\left\|\phi_{\tau}(\cdot+\tau)c(\cdot+\tau,\cdot)\right\|_{\mathrm{L}^{p}(\Omega_{0,2})}\|\theta\|_{\mathrm{L}^{p^{\prime}}(\Omega_{0,2})}.$
@@
$+C_{SOR}(\Omega,N,p^{\prime})\frac{|d_{a}-d_{c}|}{d_{a}+d_{c}}\left\|\phi_{\tau}(\cdot+\tau)\left(a+c\right)(\cdot+\tau,\cdot)\right\|_{\mathrm{L}^{p}(\Omega_{0,2})}\|\theta\|_{\mathrm{L}^{p^{\prime}}(\Omega_{0,2})}.$
@@
$\left\|\phi_{\tau}(\cdot+\tau)(a+c)(\cdot+\tau,\cdot)\right\|_{\mathrm{L}^{p}(\Omega_{0,2})}\leq\frac{K_{2}M^{\alpha}M_{1}^{\alpha}|\Omega|^{\alpha}}{\left(1-C_{SOR}(\Omega,N,p^{\prime})\frac{|d_{a}-d_{c}|}{d_{a}+d_{c}}\right)}\left\|\left(a+c\right)(\cdot+\tau,\cdot)\partial_{s}\phi_{\tau}(\cdot+\tau)\right\|_{\mathrm{L}^{p}(\Omega_{0,2})}^{1-\alpha}.$
@@
As the diffusion coefficients $d_{a}$ and $d_{c}$ satisfy the closeness condition (10), employing the $p^{\textrm{th}}$ order integrability estimate [2, Proposition 1.1, p.1186] (see Theorem 12 from the Appendix for the precise statement), we arrive at
$\beta_{0}=\|a+c\|_{\mathrm{L}^{p}(\Omega_{0,1})}\leq\left(1+\max\{d_{a},d_{c}\}\frac{|d_{c}-d_{a}|\,C^{PRC}_{\frac{d_{a}+d_{c}}{2},p^{\prime}}}{2-|d_{c}-d_{a}|\,C^{PRC}_{\frac{d_{a}+d_{c}}{2},p^{\prime}}}\right)\|a_{0}+c_{0}\|_{\mathrm{L}^{p}(\Omega)}.$
@@
$\left(b(t,x)\right)^{N}\leq 2^{N-1}\left(\left(b_{0}(x)\right)^{N}+\left(\int_{0}^{t}c(s,x)\,{\rm d}s\right)^{N}\right)\leq 2^{N-1}(1+t)^{N-1}\left(\left(b_{0}(x)\right)^{N}+\int_{0}^{t}\left(c(s,x)\right)^{N}\,{\rm d}s\right).$
@@
(26) $\left\|b(t,\cdot)\right\|_{\mathrm{L}^{N}(\Omega)}^{N}\leq 2^{N-1}(1+t)^{N-1}\Big{(}\left\|b_{0}\right\|_{\mathrm{L}^{N}(\Omega)}^{N}+\left\|c\right\|_{\mathrm{L}^{N}(\Omega_{t})}^{N}\Big{)}.$
@@
$\left\|c\right\|_{\mathrm{L}^{N}(\Omega_{\tau,\tau+1})}\leq\left\|c\right\|_{\mathrm{L}^{p}(\Omega_{\tau,\tau+1})}\,\left|\Omega\right|^{\frac{p-N}{p}}.$
@@
$\left\|b(t,\cdot)\right\|_{\mathrm{L}^{N}(\Omega)}^{N}\leq 2^{N-1}(1+t)^{N-1}\left(\left\|b_{0}\right\|_{\mathrm{L}^{\infty}(\Omega)}^{N}\,\left|\Omega\right|+C_{0}^{N}\,\left|\Omega\right|^{\frac{N(p-N)}{p}}\,(1+t)\right).$
@@
$K_{3}^{\frac{N(N-1)}{N-2}}:=2^{N-1}\left(M_{2}|\Omega|\right)^{\frac{N}{N-2}}\left(\left\|b_{0}\right\|_{\mathrm{L}^{\infty}(\Omega)}^{N}\,\left|\Omega\right|+C_{0}^{N}\,\left|\Omega\right|^{\frac{N(p-N)}{p}}\right).$
@@
$\left\|b(t,\cdot)\right\|_{\mathrm{L}^{N}(\Omega)}\leq K_{3}^{\frac{N-1}{N-2}}\left(M_{2}|\Omega|\right)^{-\frac{1}{N-2}}(1+t).$
@@
$\left\|b(t,\cdot)\right\|_{\mathrm{L}^{\frac{N}{2}}(\Omega)}\leq\left\|b(t,\cdot)\right\|_{\mathrm{L}^{N}(\Omega)}^{\frac{N-2}{N-1}}\,\left\|b(t,\cdot)\right\|_{\mathrm{L}^{1}(\Omega)}^{\frac{1}{N-1}}.$
@@
$\left\|c\right\|_{\mathrm{L}^{\frac{7}{2}}(\Omega_{\tau,\tau+1})}$ $\leq K_{c}(1+t)^{\mu_{c}}\qquad\mbox{ for }\,\tau\geq 0.$
@@
$\frac{1}{p}=\frac{j}{N}+\left(\frac{1}{r}-\frac{m}{N}\right)\alpha\ +\frac{1-\alpha}{q}\qquad\mbox{ and }\qquad\frac{j}{N}\leq\alpha\leq 1.$
@@
$\left\|\sqrt{a}(t,\cdot)\right\|_{\mathrm{L}^{\frac{10}{3}}(\Omega)}\leq K_{5}\left\|\nabla\sqrt{a}(t,\cdot)\right\|_{\mathrm{L}^{2}(\Omega)}^{\frac{3}{5}}\,\left\|\sqrt{a}(t,\cdot)\right\|_{\mathrm{L}^{2}(\Omega)}^{\frac{2}{5}}+K_{6}\left\|\sqrt{a}(t,\cdot)\right\|_{\mathrm{L}^{2}(\Omega)}.$
@@
$=2^{\frac{7}{3}}K_{5}^{\frac{10}{3}}\left\|\nabla\sqrt{a}(t,\cdot)\right\|_{\mathrm{L}^{2}(\Omega)}^{2}\,M_{1}^{\frac{4}{3}}\,\left|\Omega\right|^{\frac{4}{3}}+2^{\frac{7}{3}}K_{6}^{\frac{10}{3}}\,M_{1}^{\frac{10}{3}}\,\left|\Omega\right|^{\frac{10}{3}}.$
@@
$\leq 2^{\frac{2}{3}}(1+t)^{\frac{5}{3}}\left(\left\|b_{0}\right\|_{\mathrm{L}^{\infty}(\Omega)}^{\frac{5}{3}}\,\left|\Omega\right|+K_{7}^{\frac{5}{3}}\right).$
@@
The species $a$ is a positive subsolution of the following equation in the time interval $(0,2)$.
(30) $\partial_{t}a(t,x)-d_{a}\Delta a(t,x)\leq&c\qquad\qquad\qquad\mbox{ in }\Omega_{0,2}\\ \nabla a(t,x)\cdot n=&0\qquad\qquad\quad\ \ \mbox{ on }\partial\Omega_{0,2}\\ a(0,x)=&a_{0}\qquad\quad\qquad\ \ \mbox{ in }\Omega.\end{aligned}\right.$
The solution corresponding to the equation (34) can be expressed as:
$a\leq\int_{\Omega}G_{d_{a}}(t,0,x,y)a_{0}(y)\rm{d}y+\int_{0}^{t}\int_{\Omega}G_{d_{a}}(t,s,x,y)c(s,y)\rm{d}y\rm{d}s.$
where $G_{d_{a}}$ denotes the Green’s function associated with the operator $\partial_{t}-d_{a}\Delta$ with Neumann boundary condition. We use the fact that ${a_{0}(y)\in\mathrm{L}^{\infty}(\Omega)}$ and ${\int_{\Omega}G_{d_{a}}(t,0,x,y)\rm{d}y\leq 1}$ for all $t\in(0,2)$. It yields
$\left|\int_{\Omega}G_{d_{a}}(t,0,x,y)a_{0}(y)\rm{d}y\right|\leq\|a_{0}\|_{\mathrm{L}^{\infty}(\Omega)}.$
We use the following Green’s function estimate is from [10][3]: there exists a constant $K_{1}>0$, depending only on the domain, such that
$0\leq G_{d_{a}}(t_{1},s,x,y)\leq\frac{K_{1}}{(4\pi(t_{1}-s))^{\frac{N}{2}}}e^{-\kappa\frac{\|x-y\|^{2}}{(t_{1}-s)}}=:g_{d_{a}}(t_{1}-s,x-y),$
$\left\|g_{d}\right\|_{\mathrm{L}^{z}(\Omega_{0,2})}\leq K_{2}\qquad\forall z\in\left[1,1+\frac{N}{2}\right)$
for some constant positive $K_{2}$ depending on $z$. We, in particular choose $\frac{45}{28}$, which is an admissible number for $N=3$. As we have ${1+\frac{2}{9}=\frac{28}{45}+\frac{3}{5}}$, Young’s convolution inequality yields
$\left\|\int_{0}^{t}\int_{\Omega}G_{d_{a}}(t,s,x,y)c(s,y)\rm{d}y\rm{d}s\right\|_{\mathrm{L}^{\frac{9}{2}}(\Omega_{0,2})}\leq\|g_{d}\|_{\mathrm{L}^{\frac{45}{28}}(\Omega_{0,2})}\|c\|_{\mathrm{L}^{\frac{5}{3}}(\Omega_{0,2})}\leq K_{2}\|c\|_{\mathrm{L}^{\frac{5}{3}}(\Omega_{0,2})}.$
Exploiting the non-negativity of the species $a$ and using the estimate from (28), we obtain
$\|a\|_{\mathrm{L}^{4.5}(\Omega_{0,2})}\leq\|a_{0}\|_{\mathrm{L}^{\infty}(\Omega)}2^{\frac{2}{9}}|\Omega|^{\frac{2}{9}}+K_{2}K_{7}.$
@@
{$-[\partial_{t}\phi+d_{c}\Delta\phi]=&\Theta\qquad\ \ \ \mbox{in}\ \Omega_{0,2}\\ \nabla\phi.n=&0\qquad\mbox{on}\ (0,2)\times\partial\Omega\\ \phi(2)=&0\qquad\ \ \ \ \mbox{in}\ \Omega.\end{aligned}\right.$
We have $\phi\geq 0$, and the following estimates for a constant $C_{q,d_{c}}>0$ ( $q\in(1,\infty)$, arbitrary)[12][1]
$\|\phi_{t}\|_{\mathrm{L}^{q}(\Omega_{0,2})}+\|\Delta\phi\|_{\mathrm{L}^{q}(\Omega_{0,2})}+\sup\limits_{s\in[0,2]}\|\phi(s)\|_{\mathrm{L}^{q}(\Omega)}+\|\phi\|_{\mathrm{L}^{q}((0,2)\times\partial\Omega)}\leq C_{q,d_{c}}\|\Theta\|_{\mathrm{L}^{q}(\Omega_{0,2})}.$